# Transic: Sim-to-Real Policy Transfer by Learning from Online Correction

### v1

@@ Acknowledgments
We are grateful to Josiah Wong, Chengshu (Eric) Li, Weiyu Liu, Wenlong Huang, Stephen Tian, Sanjana Srivastava, and the SVL PAIR group for their helpful feedback and insightful discussions. This work is in part supported by the Stanford Institute for Human-Centered AI (HAI), ONR MURI N00014-22-1-2740, ONR MURI N00014-21-1-2801, and Schmidt Sciences. Ruohan Zhang is partially supported by Wu Tsai Human Performance Alliance Fellowship.
@@ A.2.3 Insert
$d=1-\tanh{\left(\frac{10}{2}d_{position}+d_{vertical}\right)}$ (A.5)
@@ A.3.1 Model Details
We use feed-forward policies [155] in RL training. It consists of MLP encoders to encode proprioceptive and privileged vector observations, and unimodal Gaussian distributions as the action head. Model hyperparameters are listed in Table A.III.
@@ A.3.2 Domain Randomization
Simulation
Gravity Additive $\mathcal{U}(0.0,0.4)$
@@ A.4.6 Training Details
To regularize point-cloud features, we separately collect a dataset containing 59 pairs of matched point clouds in simulation and reality. One pair from them is visualized in Fig A.2. Student policies are trained by minimizing Eq. 1, where we set $\beta=10^{-3}$. We use the Adam optimizer [159] with a learning rate of $10^{-4}$ during training. We periodically roll out student policies in simulation for $1,000$ episodes. We then select the checkpoint that corresponds to the highest success rate to use as the base policy in the real-world learning stage.
@@ B.1 Hardware Setup
As shown in Fig. A.3, our system consists of a Franka Emika 3 robot mounted on the tabletop. We use four fixed cameras and one wrist camera for point cloud reconstruction. They are three RealSense D435 and two RealSense D415. There is also a 3d-printed three-sided wall glued on top of the table to provide external support. We use a soft gripper for better grasping (Fig A.4). We use a joint position controller from the Deoxys library [162] to control our robot at 1000 Hz.
@@ B.2 Obtaining Point Clouds from Multi-View Cameras
Figure A.4: A soft gripper that can bend for better grasping.
@@ B.3 Human-in-the-Loop Data Collection
This data collection procedure is illustrated in Algorithm 1. As shown in Fig. A.6, we use a 3Dconnexion SpaceMouse as the teleoperation device. We design a specific UI (Fig. A.7) to facilitate the synchronized data collection. Here, the human operator will be asked to intervene or not. The operator answers through keyboard. If the operator does not intervene, the base policy’s next action will be deployed. If the operator decides to intervene, the SpaceMouse is then activated to teleoperate the robot. After the correction, the operator can exit the intervention mode by pressing one button on the SpaceMouse. We use this system and interface to collect 20, 100, 90, and 17 trajectories with correction for tasks Stabilize, Reach and Grasp, Insert, and Screw, respectively. We use 90% of them as training data and the remaining as held-out validation sets. We visualize the cumulative distribution function of human correction in Figure A.8.
@@
      $\triangleright$ deploy the base policy for one step
@@
          $\triangleright$ update observation for the next step
@@ B.4.2 Training Details
{minted} bash (…) system: need human intervention? (y/n) user: n (deploying the next action) system: need human intervention? (y/n) user: y (correction through teleopeartion) system: exiting human intervention… (…)
Figure A.7: The UI for synchronized human-in-the-loop data collection.
To train the learned gated residual policy, we first only learn the feature encoder and the action head. We then freeze the entire model and only learn the intervention head. We opt for this two-stage training since we find that training both action and intervention heads at the same time will result in sub-optimal residual action prediction. We follow the best practice for policy training [155, 98, 164], including using learning rate warm-up and cosine annealing [89]. Training hyperparameters are listed in Table A.X.
@@ Appendix C Experiment Settings and Evaluation Details
C.1 Main Experiments (Sec. 4.2)
We evaluate all methods on four tasks for 20 trials. Each trail starts with different objects and robot poses. We make our best efforts to ensure the same initial settings when evaluating different methods. Specifically, we take pictures for these 20 different initial configurations and refer to them when resetting a new trial. See Figs. A.14, A.15, A.16, A.17 for initial configurations of tasks Stabilize, Reach and Grasp, Insert, and Screw, respectively.
C.2 Experiments with Different Sim-to-Real Gaps (Sec. 4.3)
C.2.1 Experiment Setup
We elaborate on how different sim-to-real gaps are created.
@@ C.2.1 Experiment Setup
This is done by applying random jitter to 25% points from point clouds, which corresponds to adding noise in observation space $\mathcal{O}$. We test this sim-to-real gap on the task Reach and Grasp. As visualized in Fig. A.9, with probability $P=0.6$, we apply random jitter to 25% points from the point-cloud observation. The jittering noise is sampled independently from the distribution $\mathcal{N}(0,0.03)$. We clip the noise to be within the $\pm$ 0.03 range.
@@
This is done by making the joint position controller less accurate, which corresponds to mismatched action space $\mathcal{A}$. We test this gap on the task Insert. We emulate an underactuated controller through early stopping. Concretely, at every time a new joint position goal $\mathbf{q}_{goal}$ is set, we record the distance to the goal in configuration space $d_{\mathbf{q}}=\|\mathbf{q}-\mathbf{q}_{goal}\|$ and sample a factor $\Gamma\sim\mathcal{U}(0.80,0.95)$. The controller will stop reaching the desired goal once it achieves $\Gamma$ progress, i.e., stop early when $\|\mathbf{q}-\mathbf{q}_{goal}\|\leq(1-\Gamma)d_{\mathbf{q}}$. Fig. A.10 visualizes the effect.
@@
This is done by changing the robot gripper to be shorter length as demonstrated in Fig. A.11, which corresponds to discrepancy in state space $\mathcal{S}$ and transition function $\mathcal{T}$. We test this gap on the task Screw. We notice that the 9 cm length difference incurs a significant gap.
@@
This is done by changing object surfaces and increasing friction, which corresponds to different transition function $\mathcal{T}$. We test this gap on the task Stabilize. Concretely, we attach friction tapes to the square tabletop’s surface to increase friction, hence change the dynamics (Fig. A.12).
@@
As shown in Fig. A.13, this is done by replacing the table leg with a light bulb, which corresponds to change in emitting function $\Omega$. We test this gap on the task Reach and Grasp.
@@ C.2.2 Evaluation
C.3 Data Scalability Experiments (Sec. 4.4)
In Table A.XI, we show quantitative results for scalability with human correction dataset size on four tasks.
Table A.XI: Quantitative results for scalability with human correction dataset size on four tasks.
@@
C.4 Policy Robustness Experiments (Sec. 4.5)
C.4.1 Removing Cameras
We remove two cameras and only keep three. Note that this is the same number of cameras as in FurnitureBench [90]. For tasks other than Insert, we keep the wrist camera, the right front camera, and the left rear camera. For the task Insert, we keep two front cameras and the left rear camera.
C.4.2 Suboptimal Correction Data
We simulate suboptimal correction data by injecting noise into residual actions $a^{R}$. This noise is of large magnitude, which follows the normal distribution with zero mean and standard deviation corresponding to 5% of the largest residual action in the dataset.
Appendix D Additional Experiment Results
D.1 Distilling Simulation Base Policy with Diffusion Policy
@@
We experiment with learning simulation base policies (Sec. 3.1) with the Diffusion Policy [85]. Concretely, when performing action space distillation to learn student policies, we replace the Gaussian Mixture Model (GMM) action head with the Diffusion Policy. Proper data augmentation (Table A.VIII) is also applied to robustify learned policies. Hyperparameters are provided in Table A.XII.
Table A.XII: Diffusion Policy hyperparameters.
@@
The comparison between GMMs on the real robot is shown in Table. A.XIII. We highlight two findings. First, the significant domain difference between simulation and reality generally exists regardless of different policy modeling methods. Second, since the Diffusion Policy plans and executes a future trajectory, it is more vulnerable to simulation-to-reality gaps due to planning inaccuracy and the consequent compounding error. Only executing the first action from the planned trajectory and re-planning at every step may help, but the inference latency renders the real-time execution infeasible.
Table A.XIII: The real-robot performance difference between GMM and Diffusion Policy. The policy error caused by simulation-to-reality gaps will be amplified by the Diffusion Policy because it plans and executes a future trajectory.

### v2

@@ Acknowledgments
We are grateful to Josiah Wong, Chengshu (Eric) Li, Weiyu Liu, Wenlong Huang, Stephen Tian, Sanjana Srivastava, and the SVL PAIR group for their helpful feedback and insightful discussions. This work is in part supported by the Stanford Institute for Human-Centered AI (HAI), ONR MURI N00014-22-1-2740, ONR MURI N00014-21-1-2801, and Schmidt Sciences. Ruohan Zhang is partially supported by the Wu Tsai Human Performance Alliance Fellowship.
@@ A.2.3 Insert
$d=1-\tanh{\left(\frac{10}{2}\left(d_{position}+d_{vertical}\right)\right)}$ (A.5)
@@ A.3.1 Model Details
We use feed-forward policies in RL training. It consists of MLP encoders to encode proprioceptive and privileged vector observations, and unimodal Gaussian distributions as the action head. Model hyperparameters are listed in Table A.III.
@@ A.3.2 Domain Randomization
Parameter Type Distribution
@@
Parameter Type Distribution
Simulation
Gravity Additive $\mathcal{U}(0.0,0.4)$
@@ A.4.6 Training Details
To regularize point-cloud features, we separately collect a dataset containing 59 pairs of matched point clouds in simulation and reality. One pair from them is visualized in Fig A.2. Student policies are trained by minimizing the loss in Sec. 2.2, where we set $\beta=10^{-3}$. We use the Adam optimizer [112] with a learning rate of $10^{-4}$ during training. We periodically roll out student policies in simulation for $1,000$ episodes. We then select the checkpoint that corresponds to the highest success rate to use as the base policy in the real-world learning stage.
@@ B.1 Hardware Setup
As shown in Fig. A.3, our system consists of a Franka Emika 3 robot mounted on the tabletop. We use four fixed cameras and one wrist camera for point cloud reconstruction. They are three RealSense D435 and two RealSense D415. There is also a 3d-printed three-sided wall glued on top of the table to provide external support. We use a joint position controller from the Deoxys library [115] to control our robot at 1000 Hz.
@@ B.3 Human-in-the-Loop Data Collection
This data collection procedure is illustrated in Algorithm 1. As shown in Fig. A.5, we use a 3Dconnexion SpaceMouse as the teleoperation device. We design a specific UI to facilitate the synchronized data collection. Here, the human operator will be asked to intervene or not. The operator answers through keyboard. If the operator does not intervene, the base policy’s next action will be deployed. If the operator decides to intervene, the SpaceMouse is then activated to teleoperate the robot. After the correction, the operator can exit the intervention mode by pressing one button on the SpaceMouse. We use this system and interface to collect 20, 100, 90, and 17 trajectories with correction for tasks Stabilize, Reach and Grasp, Insert, and Screw, respectively. We use 90% of them as training data and the remaining as held-out validation sets. We visualize the cumulative distribution function of human correction in Fig. A.6.
@@
      $\triangleright$ deploy the base policy
@@
          $\triangleright$ update the next observation
@@ B.4.2 Training Details
To train the learned gated residual policy, we first only learn the feature encoder and the action head. We then freeze the entire model and only learn the intervention head. We opt for this two-stage training since we find that training both action and intervention heads at the same time will result in sub-optimal residual action prediction. We follow the best practice for policy training, including using learning rate warm-up and cosine annealing [117]. Training hyperparameters are listed in Table A.X.
@@ Appendix C Experiment Settings and Evaluation Details
C.1 Task Definition
As shown in Fig. 3, we quantitatively benchmark four tasks. They are fundamental skills required to assemble a square table from FurnitureBench [85]. We randomize objects’ initial poses during evaluation.
  • Stabilize: The robot pushes the square tabletop to the right corner of the wall such that it remains stable in following assembly steps.
  • Reach and Grasp: The robot reaches and grasps the table leg. It needs to properly adjust the end effector’s orientation to avoid infeasible grasping poses.
  • Insert: The robot inserts the pre-grasped table leg to the far right assembly hole of the tabletop.
  • Screw: The robot’s end-effector is initialized close to an inserted table leg and it screws the table leg clockwise into the tabletop.
C.2 Main Experiments
We evaluate all methods on four tasks for 20 trials. Each trail starts with different objects and robot poses. We make our best efforts to ensure the same initial settings when evaluating different methods. Specifically, we take pictures for these 20 different initial configurations and refer to them when resetting a new trial. See Figs. A.14, A.15, A.16, A.17 for initial configurations of tasks Stabilize, Reach and Grasp, Insert, and Screw, respectively. We follow Liu et al. [91] to label reward for IQL. Full numerical results are provided in Table A.XI.
Table A.XI: Success rates per tasks. Transic outperforms all baseline methods in all four tasks.
Tasks Transic
Direct
Transfer
DR. & Data
Aug. [52]
BC
Fine-Tune
IQL
Fine-Tune
HG-Dagger [66] IWR [67] BC [86] BC-RNN [68] IQL [69]
Stabilize 100% 10% 35% 55% 0% 65% 65% 40% 40% 5%
Reach and Grasp 95% 35% 60% 35% 0% 30% 40% 25% 0% 5%
Insert 45% 0% 15% 15% 25% 35% 40% 10% 5% 0%
Screw 85% 0% 35% 50% 65% 40% 40% 15% 25% 0%
C.3 Experiments with Different Sim-to-Real Gaps
C.3.1 Experiment Setup
We explain how different sim-to-real gaps are created.
@@ C.3.1 Experiment Setup
This is done by applying random jitter to 25% points from point clouds with probability $P=0.6$, as visualized in Fig. A.7. We test this sim-to-real gap on the task Reach and Grasp. The jittering noise is sampled independently from the distribution $\mathcal{N}(0,0.03)$. We clip the noise to be within the $\pm$ 0.03 range.
@@
This is done by making the joint position controller less accurate. We test this gap on the task Insert. We emulate an underactuated controller through early stopping. Concretely, at every time a new joint position goal $\mathbf{q}_{goal}$ is set, we record the distance to the goal in configuration space $d_{\mathbf{q}}=\|\mathbf{q}-\mathbf{q}_{goal}\|$ and sample a factor $\Gamma\sim\mathcal{U}(0.80,0.95)$. The controller will stop reaching the desired goal once it achieves $\Gamma$ progress, i.e., stop early when $\|\mathbf{q}-\mathbf{q}_{goal}\|\leq(1-\Gamma)d_{\mathbf{q}}$. Fig. A.8 visualizes the effect.
@@
This is done by changing the robot gripper to be shorter length as demonstrated in Fig. A.9. We test this gap on the task Screw. We notice that the 9 cm length difference incurs a significant gap.
@@
This is done by changing object surfaces and increasing friction. We test this gap on the task Stabilize. Concretely, we attach friction tapes to the square tabletop’s surface to increase friction, hence change the dynamics (Fig. A.10).
@@
As shown in Fig. A.11, this is done by replacing the table leg with a light bulb. We test this gap on the task Reach and Grasp.
@@ C.3.2 Evaluation
C.4 Data Scalability Experiments
In Table A.XII, we show quantitative results for scalability with human correction dataset size on four tasks.
Table A.XII: Quantitative results for scalability with human correction dataset size on four tasks.
@@ C.4 Data Scalability Experiments
C.5 Ablation Studies
C.5.1 Effects of Different Gating Mechanisms
We introduce the learned gated residual policy in Sec. 2.4 where the gating mechanism controls when to apply residual actions. To assess the quality of learned gating, we compare its performance with an actual human operator performing gating. Results are shown in Table 1 (row “w/ Human Gating”). It is evident that the learned gating mechanism only incurs negligible performance drops compared to human gating. This suggests that Transic can reliably operate in a fully autonomous setting once the gating mechanism is learned.
C.5.2 Policy Robustness
We investigate the policy robustness against 1) point cloud observations with inferior quality by removing two cameras, and 2) suboptimal correction data with noise injection. We remove two cameras and only keep three. Note that this is the same number of cameras as in FurnitureBench [85]. For tasks other than Insert, we keep the wrist camera, the right front camera, and the left rear camera. For the task Insert, we keep two front cameras and the left rear camera. We simulate suboptimal correction data by injecting noise into residual actions $a^{R}$. This noise is of large magnitude, which follows the normal distribution with zero mean and standard deviation corresponding to 5% of the largest residual action in the dataset. Results are shown in Table 1 (rows “Reduced Cameras” and “Noisy Correction”). We highlight that Transic is robust to partial point cloud inputs caused by the reduced number of cameras. We attribute this to the heavy point cloud downsampling employed during training. Fishman et al. [118] echos our finding that policies trained with downsampled synthetic point cloud inputs can generalize to partial point cloud observations obtained in the real world without the need for shape completion. Meanwhile, when the correction data used to learn residual policies are suboptimal, Transic only shows a relative decrease of 6% in the average success rate. We attribute this to the advantage of our integrated deployment—when the residual policy behaves suboptimally, the base policy could still compensate for the error in subsequent steps.
C.5.3 Consistency in Learned Visual Features
To learn consistent visual features between the simulation and reality, we propose to regularize the point cloud encoder during the distillation stage. As shown in Table 1 (row “w/o Regularization”), the performance significantly decreases without such regularization, especially for tasks that require fine-grained visual features. Without it, simulation policies would overfit to synthetic point-cloud observations and hence are not ideal for sim-to-real transfer.
C.6 Qualitative Analysis and Emergent Behaviors
We examine the distribution of the collected human correction dataset. During the human-in-the-loop data collection, the probability of intervening and correcting is reasonably low ($P_{\text{correction}}\approx 0.20$). This is consistent with our intuition that, with a good base policy, interventions are not necessary for most of the time. However, they become critical when the robot tends to behave abnormally due to unaddressed sim-to-real gaps. Moreover, as highlighted in Fig. A.6, interventions happen at different times across tasks. This fact renders heuristics-based methods [119] for deciding when to intervene difficult, and further necessitates our learned residual policy. Several representative behaviors learned by Transic are demonstrated in Fig. A.12.
(a) Error Recovery
(b) Unsticking
(c) Safety-aware Actions
(d) Failure Prevention
Figure A.12: Emergent behaviors learned by Transic. a) Error recovery. Left: The robot tries to insert the table leg but the direction is wrong; Right: Transic raises the end effector and moves to the correct insertion position. b) Unsticking. Left: The robot hovers for a while and never reaches the light bulb; Right: Transic helps the robot get unstuck and move to the bulb. c) Safety-aware actions. Left: When pushing the tabletop, the gripper is too low and bends. This might damage the robot; Right: Transic compensates for the command that causes the end effector to move too low. d) Failure prevention. Left: The light bulb will fall and break after gripper opening; Right: Transic adjusts the bulb to a stable pose to prevent failure.
Appendix D Additional Experiment Results and Discussions
D.1 Empirical Justifications for Action Space Distillation
Reasons for the proposed action space distillation are twofold. The first is mainly because an OSC is hard to sim-to-real transfer, while a joint position controller can be seamlessly transferred. As suggested in Nakanishi et al. [74], an OSC requires accurate modeling of robot parameters, such as the task-space inertia matrix and gravity compensation. System identification helps but is insufficient. Furthermore, it is often the case that given the same joint torque, the end-effector moves differently in simulation and the real world. Because an OSC uses a task-space error to compute joint torques, this will lead to large joint position deviation.
The second is for better training efficiency. As shown in Fig. A.13, it is almost impossible to directly train RL with point-cloud inputs and joint position action space. Even after 7-day training, RL still shows no sign of improvement. In contrast, Transic takes around 3 days to train on NVIDIA GeForce RTX 3090 GPUs. Therefore, the distillation is important to make the training feasible.
(a) Stabilize
(b) Reach and Grasp
(c) Insert
(d) Screw
Figure A.13: Learning curves for RL with point-cloud observations and joint position actions.
D.2 Distilling Simulation Base Policy with Diffusion Policy
@@
We experiment with learning simulation base policies (Sec. 2.2) with the Diffusion Policy [120]. Concretely, when performing action space distillation to learn student policies, we replace the Gaussian Mixture Model (GMM) action head with the Diffusion Policy. Proper data augmentation (Table A.VIII) is also applied to robustify learned policies. Hyperparameters are provided in Table A.XIII.
Table A.XIII: Diffusion Policy hyperparameters.
@@
The comparison between GMMs on the real robot is shown in Table. A.XIV. We highlight two findings. First, the significant domain difference between simulation and reality generally exists regardless of different policy parameterizations. Second, since the Diffusion Policy plans and executes a future trajectory, it is more vulnerable to simulation-to-reality gaps due to planning inaccuracy and the consequent compounding error. Only executing the first action from the planned trajectory and re-planning at every step may help, but the inference latency renders the real-time execution infeasible.
Table A.XIV: The real-robot performance difference between GMM and Diffusion Policy. The policy error caused by simulation-to-reality gaps will be amplified by the Diffusion Policy because it plans and executes a future trajectory.
@@
D.3 Gating Function Justification and Conceptual Comparison
Recall several design choices in the proposed gating mechanism: 1) takes inputs of unstructured sensory observations (point cloud); 2) conditioned on base policy’s outputs for effective prediction; 3) the intervention classifier shares the same feature encoder with the residual policy; and 4) the entire pipeline is learned end-to-end. We contrast against several mechanisms from the literature.
Table A.XV: Gating mechanism conceptual comparison.
How to decide
apply gating or not
Input
Condition on
base policy’s outputs
Shared
feature encoder
Ours End-to-end learned
Point cloud
and proprioception
Yes Yes
Residual Policy Learning [79] No gating Low-dimensional state No No
Residual RL [78] No gating Low-dimensional state No No
ThriftyDAgger [119]
Thresholded based
on neural network ensemble
Low-dimensional state No No
Runtime Monitoring [104] End-to-end learned RGB and proprioception No Yes
Although this gating function and residual policy can be merged into one, we opt not to do so for better empirical performance. In the task Screw, using the gating function with the residual policy achieves a success rate of 85%, while only using the residual policy achieves a success rate of 55%. We hypothesize that this is mainly due to the data imbalance—because intervention and correction happen with a low frequency ($P_{correction}\approx 0.20$), most residual actions are zero. If we naively learn a residual policy from them, it will be biased toward predicting near-zero actions. On the other hand, by training the residual policy only on human corrections, we offload learning whether to apply the correction to a gating function and thus reduce the negative effects of unbalanced data.
D.4 Long-Horizon Tasks Statistics
We show statistics about task length from FurnitureBench [85] in Table A.XVI.
Table A.XVI: Statistics about long-horizon tasks from FurnitureBench [85].
Number of Steps Average Human Demo Length
Lamp 594 2 Minutes
Square Table 1689 6 Minutes
Appendix E Extended Preliminaries
E.1 Intervention-Based Policy Learning
We adopt an intervention-based learning framework [66, 67, 91] where a human operator can intervene and take control during the execution of the robot base policy $\pi_{B}$. Denote the human policy as $\pi_{H}$, the following combined policy is deployed during data collection:
$\pi^{deployed}=\mathds{1}^{H}\pi^{H}+\left(1-\mathds{1}^{H}\right)\pi^{B},$ (A.7)
where $\mathds{1}^{H}$ is a binary function indicating human interventions. Introducing a trajectory distribution $q(\tau)$ that consists of two observation-action distributions generated by the robot $\rho^{B}$ and human operator $\rho^{H}$, the original RL objective leads to the maximization of a variational lower bound on logarithmic return [121, 67]:
$\mathcal{J}(\theta,q)=\mathbb{E}_{q(\tau)}\left[\log R(\tau)+\log p_{\pi_{\theta}}-\log q(\tau)\right],$ (A.8)
where $p_{\pi_{\theta}}$ is the induced trajectory distribution. While the human operator optimizes Eq. A.8 through intervention and correction, the robot learner maximizes it through
$\theta=\arg\max_{\theta\in\Theta}\mathbb{E}_{(s,a)\sim q(\tau)}\left[\log\pi_{\theta}(a|s)\right].$ (A.9)
Various intervention-based policy learning methods have been derived by weighting observation-action pairs in Eq. A.9 differently. For example, HG-Dagger [66] completely ignores robot data $\mathcal{D}^{B}$ and only trains on human data $\mathcal{D}^{H}$ that contain intervention samples. This is equivalent to $q(\tau)\propto\rho^{H}$. Intervention Weighted Regression (IWR) [67] balances the data distribution by emphasizing human intervention: $q(\tau)\propto\alpha\rho^{H}+\rho^{B}$ with $\alpha=|\mathcal{D}^{B}|/|\mathcal{D}^{H}|$. Non-intervention-based methods such as traditional behavior cloning (BC) [86] only learn on $\mathcal{D}^{H}$ with full human demonstrations instead of intervention. This effectively sets $q(\tau)\propto\rho^{H}$.
Appendix F Extended Related Work
Robot Learning via Sim-to-Real Transfer
Sim-to-Real Gaps in Manipulation Tasks
Although these gaps may not be fully bridged, traditional methods to address them include system identification [50, 18, 51, 30], domain randomization [52, 13, 53, 54, 55], real-world adaptation [56], and simulator augmentation [58, 59, 60]. However, system identification is mostly engineered on a case-by-case basis. Domain randomization suffers from the inability to identify and randomize all physical parameters. Methods with real-world adaptation, usually through meta-learning [88], incur potential safety concerns during the adaptation phase. Most of these approaches also rely on explicit and domain-specific knowledge about tasks and the simulator a priori. For instance, to perform system identification for closing the embodiment gap for a quadruped, Tan et al. [18] disassembles the physical robot and carefully calibrates parameters including size, mass, and inertia. Kim et al. [32] reports that collaborative robots, such as the commonly used Franka Emika robot, have intricate joint friction that is hard to identify and randomized in typical physics simulators. To make a simulator more akin to the real world, Chebotar et al. [39] deploys trained virtual robots multiple times to refine the distributions of simulation parameters. This procedure not only introduces a significant real-world sampling effort, but also incurs potential safety concerns due to deploying suboptimal policies. In contrast, our method leverages human intervention data to implicitly overcome the transferring problem in a domain-agnostic way and also leads to a safer deployment.
Human-in-The-Loop Robot Learning